OpenStudy (anonymous):

Simplify. Assume all variables represent positive real numbers. ^3sqrtx^6y^5

OpenStudy (anonymous):

\[\LARGE \sqrt[3]{x^6y^5}\] hint.. \[\LARGE \sqrt[a]{x^b}=\left(x\right)^{\frac ba}\] and \[\LARGE \sqrt{p\cdot q}=\sqrt p \cdot \sqrt q\]